# Friedmann equations in the Codazzi parametrization of Cotton and extended theories of gravity and the Dark Sector

### 05 Dec 2023 v1

@@ Abstract
The Friedmann equations of Cotton gravity provide a simple parametrization to reproduce, by tuning a single function, the Friedmann equations of several extensions of gravity, such as $f(R)$, modified Gauss-Bonnet $f(G)$, teleparallel $f(T)$, and more. It also includes the recently proposed Conformal Killing gravity in FRW space-times. The extensions generally have the form of a Codazzi tensor that may be associated to the dark sector. Fixing it by a suitable equation of state accomodates most of the postulated models that extend $\Lambda$CDM, as the Chevallier-Polarski-Lindler model.
Cotton gravity; Extended theories of gravity; Codazzi tensor; Conformal Killing gravity; Friedmann-Robertson-Walker spacetime
@@ I Introduction
The aim of this work is to uncover this common structure, albeit the different origins of the various cosmological models. We explicitly show this in plenty of well studied extended gravity models in FRW space-times: $f(R)$, Gauss-Bonnet $f(G)$, teleparallel $f(T)$, Einsteinian cubic $f(P)$, Conformal Killing gravity (CKT), Lovelock.
@@
In Section 5 the dark sector is fixed by requesting an EoS. It accomodates in a unified description the best known redshift dependent models, as the Chevallier-Polarski-Lindler model.
@@ III.1 $\mathbf{f(R)}$ gravity
Perhaps it is the best known extended theory of gravity. It was introduced by Buchdahl in 1970 [3] and gained popularity with the works on cosmic inflation by Starobinsky [41]. Recently $f(R)$ theories are possible candidates to explain the observed cosmic acceleration [8]. The action of $f(R)$ gravity is
@@
where $S^{(m)}$ is the matter term. With the notation $f_{R}=df/dR$, the field equations are [40]:
$f_{R}R_{kl}-\frac{f}{2}g_{kl}-[\nabla_{k}\nabla_{l}-g_{kl}\square]f_{R}=\kappa T_{kl}$ (22)
@@ III.2 $\mathbf{f(G)}$ gravity
where $f_{G}=\frac{df}{dG}$. In a FRW space-time it is a perfect fluid tensor. For the spatially flat case, the Gauss-Bonnet invariant is $G=24(\dot{H}H^{2}+H^{4})$ and the Friedmann equations of $f(G)$ gravity are expressible as (eq.5 in [36])
@@ III.3 $\mathbf{f(T)}$ gravity
In the framework of gravity theories with torsion, the “teleparallel equivalent of general relativity” is the best known one. It is widely discussed in [5] and briefly reviewed in [24]. It is based on the following action (see [24] eq. 2.5):
@@
The field equations are eqs.263 in [5] (and eq.2.6 in [24]). For the spatially flat FRW space-time, the Friedmann equations of $f(T)$ gravity are expressible as (eqs 2.9, 2.10 in [24] or eqs.267, 268 in [5]):
@@ III.4 Einsteinian Cubic and $\mathbf{f(P)}$ gravity
In [16] the cosmological applications of Einsteinian cubic gravity at early and late times were investigated. The paper [34] the authors analyze the viability of the theoretical model by considering different observational features, as cosmic chronometers data, baryon acoustic oscillations, and supernovae.
@@
Then
$-\mathscr{A}=\frac{\dot{\mathscr{B}}}{H}=12\kappa\widetilde{\beta}H^{4}\dot{H}$ (41)
From eq.(21) we get eq.(39).
@@
The Friedmann equations of Cotton gravity with the perfect fluid Codazzi tensor (13) generate the Friedmann equations of cubic Einsteinian gravity with the choice (40).
In the same paper [16] the authors proposed the extension of Einsteinian cubic gravity:
@@
The field equations are still of the type $R_{kl}-\frac{1}{2}Rg_{kl}=\kappa(T_{kl}+\widetilde{H}_{kl})$ where $\widetilde{H}_{kl}$ is quite more involved. The corresponding Friedmann equations are eqs. 26, 27 in [16]:
@@
where $f_{P}=\frac{df}{dP}$ and $P=6\widetilde{\beta}H^{4}(H^{2}+2\dot{H})$.
@@
III.5 Regularized cubic Lovelock gravity in four dimensions
In Section 3 of [7] the authors focused on the four dimensional cubic Lovelock gravity in a FRW space-time. They obtained the following Friedmann equations
@@ III.5 Regularized cubic Lovelock gravity in four dimensions
where $J^{2}=H^{2}+\frac{R^{\star}}{6a^{2}}$. We dropped their cosmological constant and the stress energy tensor is multiplied by a factor 2 to match with our notation. We can state the following:
@@
If in (21) it is $-\mathscr{B}=F(J^{2})$ where $F$ is a smooth and arbitrary function of $J^{2}=H^{2}+\frac{R^{\star}}{6a^{2}}$ then
@@
where $F_{J}=\frac{dF}{dJ^{2}}$.
Proof. From (21) we have $\frac{\kappa}{2}(p+\mu)=\frac{R^{\star}}{6a^{2}}-\dot{H}+\frac{\dot{\mathscr{B}}}{2H}$. If $-\mathscr{B}=F(J^{2})$ then ˙$\dot{\mathscr{B}}=-F_{J^{2}}(2J\dot{J})$; on the other hand $J\dot{J}=H(\dot{H}-\frac{R^{\star}}{6a^{2}})$ so that ˙$\frac{\dot{\mathscr{B}}}{2H}=-(\dot{H}-\frac{R^{\star}}{6a^{2}})F_{J^{2}}$ and the Lemma is proven. normal-$\square$
Now take $F_{J}(J^{2})=\alpha J^{2}+\beta J^{4}$ and (47) are recovered.
@@ III.6 Sussman Nájera Model in Cotton gravity
In [43] eq.16, the authors introduced the following modified Friedmann equation:
@@
with an arbitrary dimension-less function $\mathcal{K}(t)$ and a constant inverse square lenght $\gamma$. Then they computed the components of the Codazzi tensor that solves Cotton gravity. Comparison with (20) gives
@@
The Codazzi tensor reads explicitly
@@
and compares with the components in eq.17 in [43].
@@ IV Comparison with conformal Killing gravity
Shortly after, we introduced in [33] a parametrization of the theory by showing that (53) are equivalent to the Einstein’s equations in which the stress-energy tensor is augmented by a divergence-free conformal Killing tensor:
@@
In this Section we investigate the connections between Cotton and Conformal Killing gravity.
@@
So $\mathscr{C}_{kl}=K_{kl}-\frac{1}{3}Kg_{kl}$ is a Codazzi tensor with $\mathscr{C}_{r}^{r}=-\frac{1}{3}K$. From $R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+K_{kl}$ we recover the paradigm (7) with the same stress-energy tensor.
@@
Rather surprisingly, we show that any FRW space-time is equipped with a Sinyukov-like tensor.
@@
The evaluation of the gradient gives: $\nabla_{i}K_{jk}=HF^{2}(-5u_{i}g_{jk}+g_{ij}u_{k}+g_{ik}u_{j})$. Contraction with $g^{jk}$: $\nabla_{i}K=-18HF^{2}u_{i}$. It turns out that $K_{jk}$ satisfies (57), i.e. it is divergence-free Sinyukov-like. $\square$
@@
V Fixing the dark sector

### 18 Dec 2023 v2

@@ Abstract
The Friedmann equations of Cotton gravity provide a simple parametrization to reproduce, by tuning a single function, the Friedmann equations of several extensions of gravity, such as $f(R)$, modified Gauss-Bonnet $f(G)$, teleparallel $f(T)$, and more. It also includes the recently proposed Conformal Killing gravity and Mimetic gravity in FRW space-times. The extensions generally have the form of a Codazzi tensor that may be associated to the dark sector. Fixing it by a suitable equation of state accomodates most of the postulated models that extend $\Lambda$CDM, as the Chevallier-Polarski-Lindler model.
Cotton gravity; Extended theories of gravity; Codazzi tensor; Conformal Killing gravity; Mimetic gravity; Friedmann-Robertson-Walker spacetime
@@ I Introduction
The aim of this work is to uncover this common structure, albeit the different origins of the various cosmological models. We explicitly show this in plenty of well studied extended gravity models in FRW space-times: $f(R)$, Gauss-Bonnet $f(G)$, teleparallel $f(T)$, Einsteinian cubic $f(P)$, Conformal Killing gravity, Lovelock.
@@
In Section 5 we prove that the field equations of Mimetic gravity become the Cotton equations if and only if the hosting space-time is Generalized Robertson Walker, and FRW space-times are a special case.
In Section 6 the dark sector is fixed by requesting an EoS. It accomodates the best known redshift dependent models, as the Chevallier-Polarski-Lindler model.
@@ III.1 $\mathbf{f(R)}$ gravity
Perhaps it is the best known extended theory of gravity. It was introduced by Buchdahl in 1970 [4] and gained popularity with the works on cosmic inflation by Starobinsky [54]. Recently $f(R)$ theories are possible candidates to explain the observed cosmic acceleration.
Investigations to explain both dark energy and inflation were pursued in the papers by Cognola et al. [19], Nojiri and Odintsov [44, 47]. Capozziello considered $f(R)$ to discuss the issue of quintessence [8]. For general reviews on $F(R)$ see [50, 48, 53].
The action of $f(R)$ gravity is
@@
where $S^{(m)}$ is the matter term. With $f_{R}=df/dR$, the field equations are [53]:
$f_{R}R_{kl}-\frac{f}{2}g_{kl}-(\nabla_{k}\nabla_{l}-g_{kl}\square)f_{R}=\kappa T_{kl}$ (22)
@@ III.2 $\mathbf{f(G)}$ gravity
where $f_{G}=df/dG$. In a FRW space-time it is a perfect fluid tensor. For the spatially flat case, the Gauss-Bonnet invariant is $G=24(\dot{H}H^{2}+H^{4})$ and the Friedmann equations of $f(G)$ gravity are expressible as (eq.5 in [43])
@@ III.3 $\mathbf{f(T)}$ gravity
In the framework of gravity theories with torsion, the “teleparallel equivalent of general relativity” is the best known one. It is widely discussed in [6] and briefly reviewed in [31]. It is based on the action ([31] eq. 2.5):
@@
The field equations are eqs. 263 in [6] or eq. 2.6 in [31]. For the spatially flat FRW space-time, the Friedmann equations of $f(T)$ gravity are expressible as (eqs 2.9, 2.10 in [31] or eqs. 267, 268 in [6]):
@@ III.4 Einsteinian Cubic and $\mathbf{f(P)}$ gravity
In [23] the cosmological applications of Einsteinian cubic gravity at early and late times were investigated. In [41] the viability of the theoretical model is analyzed, by considering observational features such as cosmic chronometers data, baryon acoustic oscillations, and supernovae.
@@
Then $-\mathscr{A}=\dot{\mathscr{B}}/H=12\kappa\widetilde{\beta}H^{4}\dot{H}$. From eq.(21) we get eq.(39).
@@
The Friedmann equations of Cotton gravity with the perfect fluid Codazzi tensor (13) are the Friedmann equations of cubic Einsteinian gravity with the choice (40).
In the same paper [23] the authors proposed the $f(P)$ extension of Einsteinian cubic gravity:
@@
The field equations are still of the type $R_{kl}-\frac{1}{2}Rg_{kl}=\kappa(T_{kl}+\widetilde{H}_{kl})$ where $\widetilde{H}_{kl}$ is quite more involved. The Friedmann equations are eqs. 26, 27 in [23]:
@@
where $f_{P}=df/dP$ and $P=6\widetilde{\beta}H^{4}(H^{2}+2\dot{H})$.
@@
III.5 Regularized cubic Lovelock gravity
In Section 3 of [13] the authors focused on the cubic Lovelock gravity in a 4-dimensional FRW space-time. They obtained the following Friedmann equations
@@ III.5 Regularized cubic Lovelock gravity
where $J^{2}=H^{2}+\frac{R^{\star}}{6a^{2}}$. We dropped their cosmological constant and the stress energy tensor is multiplied by a factor 2 to match our notation. We state the following:
@@
If in (21) we put $-\mathscr{B}=F(J^{2})$ where $F$ is a smooth arbitrary function of $J^{2}=H^{2}+\frac{R^{\star}}{6a^{2}}$ then
@@
where $F_{J}=dF/dJ^{2}$.
Proof. From (21) we have $\frac{\kappa}{2}(p+\mu)=\frac{R^{\star}}{6a^{2}}-\dot{H}+\frac{\dot{\mathscr{B}}}{2H}$. If $-\mathscr{B}=F(J^{2})$ then ˙$\dot{\mathscr{B}}=-2F_{J^{2}}J\dot{J}$. On the other hand $J\dot{J}=H(\dot{H}-\frac{R^{\star}}{6a^{2}})$ so that ˙$\frac{\dot{\mathscr{B}}}{2H}=-(\dot{H}-\frac{R^{\star}}{6a^{2}})F_{J^{2}}$ and the Lemma is proven. normal-$\square$
Choose $F_{J}(J^{2})=\alpha J^{2}+\beta J^{4}$ and (45)(46) are recovered.
@@ III.6 Sussman Nájera Model in Cotton gravity
In [56, 57] the authors introduced the following modified Friedmann equation:
@@
with an arbitrary dimension-less function $\mathcal{K}(t)$ and a constant $\gamma$. Then they computed the components of the Codazzi tensor that solves Cotton gravity. Comparison of (48) with (20) gives $\mathscr{B}$ whence $\mathscr{A}$ is computed:
@@
The expression of the Cotton tensor
@@
compares with the components in eq.19 evaluated in [57].
@@ IV Comparison with conformal Killing gravity
Shortly after in [40] we introduced a parametrization of the theory by showing that (52) is equivalent to the Einstein’s equation in which the stress-energy tensor is augmented by a divergence-free conformal Killing tensor:
@@
Here we investigate the connections between Cotton and Conformal Killing gravity.
@@
Then $\mathscr{C}_{kl}=K_{kl}-\frac{1}{3}Kg_{kl}$ is a Codazzi tensor with $\mathscr{C}_{r}^{r}=-\frac{1}{3}K$. From $R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+K_{kl}$ we recover the paradigm (7) with the same stress-energy tensor.
@@
We show that, rather surprisingly, any FRW space-time is equipped with a Sinyukov-like tensor.
@@
Next evaluate $\nabla_{i}K_{jk}=HF^{2}(-5u_{i}g_{jk}+g_{ij}u_{k}+g_{ik}u_{j})$. Contraction with $g^{jk}$: $\nabla_{i}K=-18HF^{2}u_{i}$. It turns out that $K_{jk}$ satisfies (56), i.e. it is divergence-free Sinyukov-like. $\square$
@@
V Comparison with Mimetic gravity
In 2013, Chamseddine and Mukhanov [14, 15, 16] proposed a modification of GR where the conformal degree of freedom is distinguished. This is done by parametrizing the physical metric tensor $g_{kl}$ in terms of an auxiliary metric $\widetilde{g}_{kl}$ and a scalar field $\phi$, called mimetic field:
$g_{kl}(\widetilde{g},\phi)=-(\widetilde{g}^{pq}\nabla_{p}\phi\nabla_{q}\phi)\,\widetilde{g}_{kl}$ (63)
where $\widetilde{g}^{pq}\equiv(\widetilde{g}^{-1})_{pq}$. Then $g^{kl}=-(\widetilde{g}^{pq}\nabla_{p}\phi\nabla_{q}\phi)^{-1}\,\widetilde{g}^{kl}$. The compatibility condition follows:
$g^{kl}\nabla_{k}\phi\nabla_{l}\phi=-1$ (64)
A conformal transformation of the auxiliary metric $\widetilde{g}^{\prime}_{kl}=\Omega^{2}\widetilde{g}_{kl}$ leaves the physical metric invariant. Mimetic gravity may be viewed as a conformal extension of Einstein theory, which is locally Weyl invariant: this fact was pointed out by Barvinsky [3].
The gravitational action depends upon the auxiliary metric and the mimetic field. Alternatively, it depends on the physical metric but with the constraint (64):
$S=\int d^{4}x\sqrt{-g}\left[R+\zeta(g^{pq}\nabla_{p}\phi\nabla_{q}\phi+1)-V(\phi)\right]+S^{(m)}$ (65)
$V(\phi)$ is a potential and $\zeta$ is a Lagrange multiplier. For a thorough review of Mimetic gravity see [52].
The first field equation is obtained by minimising with respect to the metric:
$R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+2\zeta\nabla_{k}\phi\nabla_{l}\phi+g_{kl}V(\phi)$ (66)
It has the form of an extended theory with dark sector explicitly represented by the mimetic field (whence the name of ‘mimetic dark matter’ in the literature).
The trace and the constraint give $2\zeta=R+T+4V$. The covariant divergence of (66) is
$02[\nabla^{k}\zeta\nabla_{k}\phi+\zeta\nabla^{k}\nabla_{k}\phi]\nabla_{l}\phi+\nabla_{l}V=$ 0 (67)
where we used $\nabla_{k}R^{k}{}_{l}-\frac{1}{2}\nabla_{l}R=0$, $\nabla^{k}T_{kl}=0$ and $\nabla_{j}(\nabla^{p}\phi\nabla_{p}\phi)=0$.
Variation of the action with respect to the mimetic field gives
$2\nabla^{p}(\zeta\nabla_{p}\phi)=-\frac{\partial V}{\partial\phi}$ (68)
Since $g^{pq}\nabla_{p}\phi\nabla_{q}\phi=-1$ the vector field $u_{k}=-\nabla_{k}\phi$ is unit time-like and closed, i.e. $\nabla_{j}$$u_{k}=\nabla_{k}u_{j}$. Then it is vorticity-free and acceleration-free:
$\nabla_{j}u_{k}=H(g_{jk}+u_{j}u_{k})+\sigma_{jk}$ (69)
being $\sigma_{jk}$ the shear tensor. The corresponding metric is (see [16, 14] and [20])
$ds^{2}=-dt^{2}+g_{\mu\nu}^{\star}({\bf x},t)dx^{\mu}dx^{\nu}$ (70)
By fixing the hypersurfaces of constant time of (70) to be of constant $\phi$, the solution of the constraint $g^{pq}\nabla_{p}\phi\nabla_{q}\phi=-1$ may be written (see [14] and reference therein or [16]):
$\phi=\pm t+const.$ (71)
Thereby choosing $\phi=t$ and using $u_{k}=-\nabla_{k}\phi$, it is $u_{0}=-1,u_{\mu}=0$.
We then conclude that the general metric for mimetic gravity is (70). In this context $V=V(t)$, while in general $\zeta$ is a function of ${\bf x}$ and $t$.
The field equations take the form
$R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+2\zeta u_{k}u_{l}+Vg_{kl}$ (72)
and (67) becomes $\nabla_{l}V=-2(\dot{\zeta}+3H\zeta)u_{l}$. Transvecting it with $u^{l}$ gives the interesting relation
$\dot{V}=2\dot{\zeta}+6H\zeta$ (73)
where we used $\nabla_{p}u^{p}=3H$ derived from (69).
Now note that (72) may be rewritten as in Cotton gravity $R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+\mathscr{C}_{kl}-g_{kl}\mathscr{C}_{r}^{r}$, with
$\mathscr{C}_{kl}=2\zeta u_{k}u_{l}+\frac{1}{3}g_{kl}(2\zeta-V)$ (74)
It is a perfect fluid tensor with $\mathscr{A}=2\zeta$ and $3\mathscr{B}=2\zeta-V$ but in general it is not Codazzi. Nevertheless, in view of (73), it is always
˙$-\frac{\dot{\mathscr{B}}}{H}=\frac{1}{3H}(\dot{V}-2\dot{\zeta})=2\zeta=\mathscr{A}$
We report Theorem 2.1 in [38] restricted to the case of vanishing acceleration:
The perfect fluid tensor $\mathscr{C}_{kl}=\mathscr{A}u_{k}u_{l}+\mathscr{B}g_{kl}$ is Codazzi if and only if: 1) $\nabla_{j}u_{k}=H(g_{jk}+u_{j}u_{k})$, 2) $\nabla_{j}H=-\dot{H}u_{j}$, 3) $\nabla_{j}\mathscr{A}=-\dot{\mathscr{A}}u_{j}$ and $\nabla_{j}\mathscr{B}=-\dot{\mathscr{B}}u_{j}$, 4) $H=-\dot{\mathscr{B}}/\mathscr{A}$.
This can be rephrased as follows:
Proposition 10
The field equation (66) of Mimetic gravity is the field equation of Cotton gravity if and only if the space-time is GRW, $V=V(t)$ and $\zeta=\zeta(t)$.
In particular, in a FRW space-time the field equations (66) are the Cotton equations.
Cotton gravity can include other versions of mimetic gravity. As noted in [15] (see also the review [13]) in order to have viable cosmological perturbations the action (65) has to include higher derivative terms. For example it is possible to add $\frac{1}{2}\gamma(\square\phi)^{2}$ being $\gamma$ a constant. The new field equations are (eq. 110 in [13]):
$R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+g_{kl}[V(\phi)+\gamma\nabla_{p}\chi\nabla^{p}\phi]$
$\qquad\qquad+2\zeta\nabla_{k}\phi\nabla_{l}\phi-\gamma[\nabla_{k}\phi\nabla_{l}\chi+\nabla_{k}\chi\nabla_{l}\phi]$ (75)
where $\square\phi=\chi$. The background is a FRW space-time. Since $u_{k}=-\nabla_{k}\phi$ we get $3H=-\square\phi$. Moreover, recalling that $\nabla_{k}H=-\dot{H}u_{k}$, it is $\nabla_{k}\chi=-3\nabla_{k}H=3\dot{H}u_{k}$ and the previous equation rewrites as
$R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+$ $2(\zeta+3\gamma\dot{H})u_{k}u_{l}$
$+g_{kl}[V(\phi)+3\gamma\dot{H}]$ (76)
Thus we recognize
$\mathscr{C}_{kl}=2(\zeta+3\gamma\dot{H})u_{k}u_{l}+\frac{1}{3}g_{kl}(2\zeta-V+3\gamma\dot{H})$ (77)
This is again perfect fluid, with $\mathscr{A}=2(\zeta+3\gamma\dot{H})$ and $3\mathscr{B}=2\zeta-V+3\gamma\dot{H}$ being in this context $V=V(t)$ and $\zeta=\zeta(t)$. The covariant divergence of (76) gives the conservation law
$02[3H\zeta+9\gamma H\dot{H}+\dot{\zeta}+3\gamma\ddot{H}]u_{l}+\nabla_{l}V+3\gamma\nabla_{l}\dot{H}=$ 0 (78)
Transvecting this with $u^{l}$ gives
$\gamma\ddot{H}=2H(\zeta+3\gamma\dot{H})-\frac{2\dot{\zeta}}{3}+\frac{\dot{V}}{3}$ (79)
Thus $3\dot{\mathscr{B}}=2\dot{\zeta}-\dot{V}+3\gamma\ddot{H}$ and using (79) it is ˙$\frac{\dot{\mathscr{B}}}{H}=2(\zeta+3\gamma\dot{H})=-\mathscr{A}$ and (77) is a Codazzi tensor.
We have proven the following
Proposition 11
In a FRW space-time the field equations (76) are the Cotton equations.
A more general Lagrangian containing a term proportional to $(\nabla_{k}\nabla_{l}\phi)^{2}$ was investigated in [12]. In [49] the authors introduced Mimetic $f(R)$ gravity.
VI Fixing the dark sector